OpenStudy (anonymous):

Need Help ASAP! Frank is in an airplane flying at 14,080 ft above sea level. He wants to land the airplane on a runway that is 22 ft below sea level. What was the total vertical distance between his current height and the runway?

OpenStudy (chosenmatt):

what is 14,080 plus 22ft?

OpenStudy (anonymous):

14,102

OpenStudy (chosenmatt):

lol im praticing not giving direct answers lol

OpenStudy (anonymous):

OH!!!! SORRY!

OpenStudy (chosenmatt):

no problem

OpenStudy (anonymous):

There's another one >.<.The variable t represents the temperature at 4 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Mansfield dropped 17F between 4 p.m. and 11 p.m. The temperature at 11 p.m. was –11F. What was the temperature at 4 p.m.? A. 17 + 11 = t B. 17 – t = –11 C.11 – 17 = t D. t – 17 = –11

OpenStudy (anonymous):

nvm i got it